On average, 100 guests stay in a hotel on any given day. The hotel turns over its inventory of guests 7 times a week. On an average day, how many guests check into the hotel?
Answer #1

WIP = 100 guests
Number of inventory turnovers in a week = 7

So, Flow time = 1/7 weeks = 1 day (assuming 1 week = 7 days)

WIP = Flow rate x Flow time (Little's law)

or, Flow rate = WIP / Flow time

or, Flow rate = 100 / 1 = 100 guests per day

So, on an average, 100 guests check into the hotel per day.